OpenStudy (anonymous):

Select one of the factors of 3x2 + 4x – 4 (x + 4) (3x + 2) (3x – 2) (x – 2)

OpenStudy (callisto):

Can you try to do the factorization first? Hint , 4 can be expressed in 2x2

OpenStudy (anonymous):

hint.. use quadratic formula ... \[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (callisto):

 So the 2 factors are (3x-2)(x+2) Choose your own answer please